### v1

# CITYWALK: Enhancing LLM-Based C++ Unit Test Generation via Project-Dependency Awareness and Language-Specific Knowledge

Abstract.
Unit testing plays a pivotal role in the software development lifecycle, as it ensures code quality. However, writing high-quality unit tests remains a time-consuming task for developers in practice. More recently, the application of large language models (LLMs) in automated unit test generation has demonstrated promising results. Existing approaches primarily focus on interpreted programming languages (e.g., Java), while mature solutions tailored to compiled programming languages like C++ are yet to be explored. The intricate language features of C++, such as pointers, templates, and virtual functions, pose particular challenges for LLMs in generating both executable and high-coverage unit tests. To tackle the aforementioned problems, this paper introduces CITYWALK, a novel LLM-based framework for C++ unit test generation. CITYWALK enhances LLMs by providing a comprehensive understanding of the dependency relationships within the project under test via program analysis. Furthermore, CITYWALK incorporates language-specific knowledge about C++ derived from project documentation and empirical observations, significantly improving the correctness of the LLM-generated unit tests. We implement CITYWALK by employing the widely popular LLM GPT-4o. The experimental results show that CITYWALK outperforms current state-of-the-art approaches on a collection of eight popular C++ projects. Our findings demonstrate the effectiveness of CITYWALK in generating high-quality C++ unit tests.
Unit Test Generation, Large Language Model, Program Analysis, Language-Specific Knowledge, Retrieval-Augmented Generation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software testing and debugging
1. Introduction
The C++ programming language, widely renowned for its efficiency, scalability, and security, plays a crucial role in developing basic software such as operating systems and compilers. As these systems increase in size, however, ensuring code reliability becomes increasingly challenging in the fast-evolving software development process (Othmane et al., 2014). Unit testing (Runeson, 2006; Daka and Fraser, 2014) serves as a fundamental technique in this pursuit, providing a robust means of validating individual software units in isolation from the rest of the system. By independently verifying each unit, developers can detect and rectify defects early in the software development lifecycle, thereby improving overall code quality. Nevertheless, writing high-quality C++ unit test cases becomes a challenging and time-consuming task for developers in practice when the complexity of software systems grows (Garousi and Zhi, 2013). More recently, the application of large language models (LLMs) in unit test case generation has been extensively explored in both academia and industry (Schäfer et al., 2024; Yuan et al., 2024; Tang et al., 2024; Alshahwan et al., 2024b, a). For instance, Yuan et al. (Yuan et al., 2024) conducted a comprehensive evaluation of using ChatGPT in automatically generating unit test cases for Java projects via both quantitative analysis and user studies. Their findings indicate that the ChatGPT-generated test cases exhibit commendable readability, thereby affirming the feasibility of an LLM-based technological approach.
Figure 1. Limitations of LLMs in Generating C++ Unit Test Cases.
[…]
Although LLM-based approaches have achieved remarkable performance in unit test case generation, the test code generated by LLMs still encounters issues, including compilation errors and assertion failures. Furthermore, current state-of-the-art approaches and tools (Sapozhnikov et al., 2024; Ryan et al., 2024; Chen et al., 2024a; Yuan et al., 2024) primarily concentrate on interpreted programming languages such as Java and Python, with limited mature solutions yet available for generating unit test cases tailored to compiled programming language like C++. Particularly, the intricate language features of C++ present substantial challenges in utilizing LLMs to generate executable unit test cases with high coverage. To underscore the limitations of LLMs in the generation of C++ unit test cases, Figure 1 illustrates three motivating examples that demonstrate typical scenarios in which the most advanced LLM GPT-4o, when provided with the basic prompt  [^1] Following ChatTester (Yuan et al., 2024), the basic prompt comprises the task description for unit test generation, a requirement to understand program’s intent, the source code of the focal method, and the dependency contexts within the focal class file., fails to generate correct unit test cases for the corresponding focal methods (i.e., method under test) in the open-source C++ projects.
  • Limitation 1: Missing code-agnostic contexts pertaining to the project configuration dependencies for LLMs to generate compilable code. Existing LLM-based approaches (Chen et al., 2024a; Yuan et al., 2024) leverage static analysis techniques to extract code dependency contexts related to the focal method from the focal class file, which aims to ensure the syntactic correctness of LLM-generated test code. However, LLMs may also produce compilation errors due to a lack of awareness regarding dependencies specified within the project’s configuration. As shown in the Error Message output by Failed Test Case ➊, the test code generated by GPT-4o fails to compile because the configuration file for the tinyxml2 project does not include the usage of the third-party library gtest. Given the absence of such code-agnostic contexts (e.g., the availability of specific programming framework) in the provided basic prompt, there is a high likelihood that GPT-4o will default to directly using the GoogleTest framework to generate C++ unit test cases for tinyxml2.
  • Limitation 2: Lack of cross-file intended behavior in the corresponding project as guidance for LLMs to generate correct assertions. Considering the example of Focal Method (b), the functionality of the Translate method in the tag class file is to invoke the relevant interface methods provided by the Directives class to translate various YAML tags in accordance with their respective types. In this instance, the variable type is defined within the header file of tag, while the contextual information related to the invoked interface methods resides in the class file of Directives. Given that these cross-file dependency contexts within the yaml-cpp project are not encompassed in the basic prompt, comprehending the functional intent of Translate alone does not sufficiently assist GPT-4o in grasping the usage of Translate. Consequently, Failed Test Case ➋ is unable to configure the correct YAML tag type to align with the target translation pattern. This misalignment results in the actual outcome that deviates from the expected result, thereby leading to an assertion failure in the generated test code.
  • Limitation 3: LLMs struggle to correctly generate complex test code without an understanding of C++ language-specific domain knowledge. In practice, the focal method may need to interact with complex data objects defined in dependent class files. In such cases, developers typically create virtual objects using mocking techniques (Saff and Ernst, 2004; Thomas and Hunt, 2002) to simulate the behavior of real external dependencies. However, to effectively generate such complex test code, LLMs must grasp the principles and applicable scenarios related to mocking (Spadini et al., 2019, 2017). In Focal Method (c), the decode method is required to return a boolean value by verifying whether node is a scalar. As illustrated in the generated Failed Test Case ➌, GPT-4o attempts to simulate a Node class object to mimic the test input. The Error Message indicates that the methods invoked on node are non-virtual functions. As a consequence, GPT-4o generates test code that contains compilation errors, stemming from an insufficient understanding of such complex language feature of C++ mocking.
To tackle the aforementioned limitations, we propose CITYWALK, a novel framework designed to enhance the capabilities of LLMs in generating high-quality C++ unIt Test cases by providing project-dependencY aWAreness and Language-specific Knowledge. Our key insight lies in enabling LLMs to act as skilled developers, leveraging specialized knowledge and global comprehension of project dependencies to perform unit testing effectively. The main ideas of CITYWALK are outlined as follows:
  • Empowering LLMs with an awareness of project-level dependency relationships through program analysis. The project dependencies employed by CITYWALK are categorized into two types: cross-file data dependencies and configuration dependencies. To conduct cross-file data dependency analysis, CITYWALK initially performs static analysis to identify the header and source code files within the project that exhibit a dependency-chain relationship with the focal class file. Subsequently, CITYWALK utilizes the abstract syntax tree (AST) to extract the relevant cross-file data dependency contexts associated with the focal method from these identified files. Furthermore, our observations indicate that when the project under test necessitates specific versions of the compilation environment or third-party libraries, the probability of encountering compilation errors in the LLM-generated unit tests increases substantially. By analyzing the configuration files of the project, CITYWALK gathers critical information regarding the third-party library usage and the requirements of the compilation environment. These configuration dependencies are explicitly provided to LLMs to reduce the occurrence of compilation errors.
  • Augmenting LLMs with intention-guiding information related to the focal method via a hybrid retrieval strategy. Software repositories often contain extensive documentation, such as requirements specifications and API references (Maalej and Robillard, 2013). Utilizing this natural language information can further aid LLMs in understanding the functional logic of the focal method. Moreover, when the focal method has long dependency chains or complex initialization of dependent objects, incorporating relevant code snippets, especially those that invoke the focal method or demonstrate its initialization, into the prompts can greatly help guide LLMs in understanding the real-world usage patterns of the focal method, thereby compensating for the limitations of static analysis techniques. To facilitate the retrieval of both natural language documentation and programming language code snippets, CITYWALK employs a hybrid strategy designed to efficiently extract intention-guiding information from bimodal sources, ensuring an accurate and contextually relevant retrieval process.
  • Prompting LLMs with step-by-step instructions and language-specific knowledge derived from empirical observations. CITYWALK decomposes the unit test generation task into smaller, more specific subtasks, structured as step-by-step instructions. The goal is to streamline the generation of C++ unit test cases by providing LLMs with detailed procedural guidance. Within CITYWALK, LLMs are utilized to: (1) infer both the intention and dependencies of the focal method; (2) generate a test suite for the focal method by leveraging the provided guidance; and (3) refine the generated test cases using language-specific knowledge. To perform Step (3), we further conduct an empirical analysis of the LLM-generated failed test cases for building language-specific domain knowledge tailored to C++. This knowledge-driven approach assists in guiding LLMs to emulate experienced developers in generating accurate test cases.
We implement a prototype of CITYWALK using GPT-4o and evaluate it on a collected benchmark comprising 1057 focal methods from eight C++ projects. We conduct a comparative analysis of CITYWALK against six state-of-the-art baselines. Our evaluation demonstrates that CITYWALK outperforms all baselines in both compilation success rate and code coverage. We further integrate CITYWALK as a plugin within the open-source code LLMs, with the aim of proving its generalization capabilities and effectiveness in C++ unit test generation. Our contributions can be summarized as follows:
  • We present the first attempt at enhancing the capabilities of LLMs for C++ unit test generation by leveraging project-dependency awareness and language-specific knowledge, enabling LLMs to function like human developers in writing correct unit test cases.
  • We thoroughly evaluate CITYWALK on a diverse collection of open-source projects, thereby substantiating the effectiveness of each component within CITYWALK. Furthermore, we illustrate its capacity to generalize in generating high-quality unit tests across various LLMs.
  • We publicly release the artifacts (Zhang, 2024) of CITYWALK on Zenodo to facilitate the reproduction of CITYWALK by the researchers in the software engineering community.
Article Organization. The remainder of this paper is organized as follows: Section 2 introduces in detail the proposed framework. Section 3 and Section 4 provide the experimental setups and results of our research. Section 5 presents case studies and discloses the threats to the validity of our approach. Section 6 describes the related work. Section 7 draws conclusions of our research.
2. Methodology
In this paper, we introduce CITYWALK, a novel LLM-based framework for C++ unit test generation that addresses the limitations outlined in Section 1. As illustrated in Figure 2, CITYWALK consists of five stages designed to enhance LLMs in generating high-quality C++ unit test cases. It effectively integrates the strengths of program analysis with retrieval-augmented strategies, empowering LLMs with developer-like capabilities through three key aspects of guidance.
Figure 2. Overview of CITYWALK.
[…]
2.1. Task Formulation
We first provide a formal definition of the unit test generation task. Let $\mathtt{Repo=\{SC,C,D\}}$ be the project repository, which includes a set of source code files $\mathtt{SC=\{sc_{1},sc_{2},...sc_{i}\}}$, configuration files $\mathtt{C=\{c_{1},c_{2},...c_{j}\}}$, and documentation files $\mathtt{D=\{d_{1},d_{2},...d_{k}\}}$. The methods within $\mathtt{Repo}$ are denoted as $\mathtt{M=\bigcup_{sc\in SC}M(sc)}$, where $\mathtt{M(sc)=\{m_{1},m_{2},...,m_{n}\}}$ is the set of methods in $\mathtt{sc}$. Given a focal method $\mathtt{m_{focal}\in M(sc_{focal})}$ associated with the focal class file $\mathtt{sc_{focal}}$, the objective of CITYWALK is to generate unit test case(s) $\mathtt{TC}$ for $\mathtt{m_{focal}}$ through the following stages:
➀ Project Repository Pre-Processing::
This stage involves navigating through all necessary files in $\mathtt{Repo}$, segmenting the selected files into manageable chunks, and storing the vectorized chunks in the knowledge bases $\mathtt{KB_{code}}$ and $\mathtt{KB_{docs}}$.
➁ Project Dependency Extraction::
In this stage, we identify configuration dependencies within $\mathtt{C}$ and cross-file data dependencies between $\mathtt{sc_{focal}}$ and $\mathtt{SC}$ via program analysis.
➂ Intention Context Retrieval::
This stage focuses on retrieving relevant code snippets from $\mathtt{KB_{code}}$ and documentation from $\mathtt{KB_{docs}}$ as intention contexts utilizing a hybrid strategy.
➃ Empirical Observation::
In this stage, we analyze the LLM-generated failed test cases and summarize the language-specific knowledge derived from empirical observations.
➄ Unit Test Generation::
This stage involves prompting the LLM with step-by-step instructions to generate test cases $\mathtt{TC\leftarrow LLM(PROMPT)}$, where the content of $\mathtt{PROMPT}$ includes the basic prompt, supplemented with additional guidance from Stage -, and incorporates rule-based methods to fix test cases with syntactic and compilation errors.
2.2. Project Repository Pre-Processing
In the context of program analysis-based and retrieval-augmented unit test generation, this stage involves two preliminary processes for pre-processing the given focal method and corresponding project. These processes are essential to help LLMs understand the code structure and its dependencies, thereby facilitating precise analysis and effective retrieval to support test case generation.
2.2.1. Structured Focal Context Extraction
Given the input focal class file, CITYWALK begins by parsing the source code into an AST using Clang  [^2] https://clang.llvm.org. For each focal method, CITYWALK then gathers the essential focal contexts in a structured format, providing the foundational information necessary for project dependency analysis. Figure 3 illustrates the structured focal contexts for decode in the convert class, preserving the complete code implementation of the focal method, along with imports from C++ standard libraries, third-party libraries, user-defined header files, namespace declarations, and signatures of other methods within the focal class.
Figure 3. Illustrative Example of Structured Focal Context for the decode Method in the convert Class.
[…]
2.2.2. Knowledge Base Construction
Retrieval-augmented generation (RAG) allows LLMs to leverage information from external knowledge bases for reducing hallucinations (Chen et al., 2024b). To enhance the LLM’s understanding of the focal method during unit test generation, CITYWALK constructs knowledge bases using the documentation and source code files from the project repository. CITYWALK first navigates through all representative files within the repository, including project documentation files (e.g., requirement specifications) written in natural language and source code files in programming languages. Next, CITYWALK applies text segmentation strategies to slice the documentation and code files. For instance, markdown files can be segmented into text chunks based on title levels, while source code files are divided into method-level code chunks. Additionally, natural language comments within the code are extracted separately as text chunks. To facilitate efficient semantic retrieval and similarity calculations, CITYWALK utilizes an embedding model BGE (Xiao et al., 2024) for the vectorization of segmented text chunks, storing the resulting vectors in $\mathtt{KB_{docs}}$. The segmented code chunks are directly stored in $\mathtt{KB_{code}}$ for exact match retrieval without vectorization.
2.3. Project Dependency Extraction
LLM-generated test code often fails to compile, typically due to missing dependencies or specific version requirements of the compilation environment. This limitation compels LLMs to make assumptions about the usage of undefined variables or methods based on their reasoning (Zhang et al., 2024a). To mitigate compilation errors, CITYWALK enables LLMs to be aware of relevant dependency contexts at the project level through program analysis. The project dependencies extracted by CITYWALK are classified into the following two categories:
  • Configuration Dependencies: As illustrated in Figure 4(a), CITYWALK identifies both the usage of third-party libraries and the compilation environment requirements of the project under test. CITYWALK first parses the CMakeLists files in the root and test dictionaries using regular expressions to extract configuration dependencies based on keyword-matching, including the command like target_link_libraries. This process captures details about dependent third-party libraries and their versions (e.g., gtest-1.11.0) as well as the compilation environment requirement (i.e., C++ 11) specified through set_target_properties. The parsed dependencies (framed by the green rectangle) are stored in a structured format that facilitates seamless integration with LLM prompts. This enables LLMs to understand the necessary code-agnostic dependencies of the project under test, which is essential for generating compilable test code.
[…]
Figure 4. Illustrative Example of Project Dependency Extraction.
[…]
2.4. Intention Context Retrieval
To better guide LLMs in understanding the program intention of the focal method, CITYWALK employs RAG to generate effective unit test cases by leveraging the project documentation and source code stored in the knowledge bases described in Section 2.2.2. The natural language documentation provides functionality requirements of the focal method as articulated by developers, while the source code snippets within the project offer LLMs real-world examples of focal method invocation and initialization. This approach mitigates issues related to focal methods that have dependency chains longer than two layers or are complex to initialize. Specifically, CITYWALK employs a hybrid retrieval strategy that conducts knowledge queries in two stages. For $\mathtt{KB_{docs}}$, the process begins by generating a corresponding query statement (as shown in Figure 2) based on the name of the focal method and the class to which it belongs. The embedding model BGE is then utilized to convert this query into a vector representation, capturing its semantic information and enabling alignment with the content of $\mathtt{KB_{docs}}$ in vector space. Using the retrieval algorithms provided by the vector database Faiss (Douze et al., 2024), the similarity between the query vector and the vectors in $\mathtt{KB_{docs}}$ is computed, employing cosine similarity to identify the vectors that are most similar to the query vector. Based on the similarity scores, the top-2 most relevant responses are selected as retrieval results. This setup ensures retrieval accuracy while providing sufficiently detailed guidance information, assisting in the generation of more comprehensive test cases. Preliminary practices indicate that semantic retrieval for code snippet examples often yields imprecise results (Bajracharya et al., 2010; Yan et al., 2020). Thus, CITYWALK optimally employs an exact match retrieval strategy for $\mathtt{KB_{code}}$, using regular expressions to identify examples of method invocations or initializations based on the name of the focal method.
2.5. Empirical Observation
In addition to the extracted project dependencies and retrieved intention contexts, CITYWALK further provides error patterns, along with their corresponding solution guidelines, as language-specific domain knowledge. This knowledge tailored to C++ programming language is used to prompt LLMs with instructions aimed at mitigating common errors in the task of unit test generation. To achieve this, we conduct an empirical study to evaluate the correctness of C++ test cases generated by existing state-of-the-art LLMs through quantitative analysis. The procedure for our empirical investigation is outlined as follows:
  • Empirical Setups: (1) Benchmark. We adopt a diverse set of real-world open-source C++ projects collected from GitHub, with the selection criteria and detailed statistics provided in Section 3.2. (2) Subject LLMs. We select two state-of-the-art LLMs to evaluate their capabilities in generating high-quality C++ unit test cases based on our collected benchmark: the open-source LLM DeepSeek-V2.5 (DeepSeek-AI, 2024a, b) and the closed-source LLM GPT-4o (OpenAI, 2023). (3) Prompt Design. As suggested by Yuan et al. (Yuan et al., 2024), we design our prompt by closely following common practices in recent unit test generation research. The prompt consists of a natural language description that explains the task to the LLM, along with the code context, which includes the focal method and other relevant code.
  • Experimental Procedure: We query the selected LLMs using our designed prompt for each focal method within the collected benchmark and consider the test cases generated by the LLMs as the output. The official API of each LLM is used with the configuration set to generate the top-1 chat completion choice and a sampling temperature of 0. The generated test cases are then placed in the test directory of the project under test, where we attempt to compile and execute them for subsequent analysis. We evaluate the correctness of the generated test cases from three perspectives. First, we use the Clang parser as a syntax checker to verify the syntactic correctness of the generated test cases. Next, we measure their compilation and execution correctness by determining whether the test cases could compile successfully and pass execution. In addition, we employ llvm-cov  [^3] https://llvm.org/docs/CommandGuide/llvm-cov.html to collect both line and branch coverage, providing an assessment of the sufficiency of the generated test cases.
  • Error Analysis: To better understand the limitations of existing LLMs in C++ unit test generation, we analyze common error patterns in the failed test cases generated by DeepSeek-V2.5 and GPT-4o. Specifically, we automatically build the error taxonomies based on the error messages encountered during compilation and execution, and then manually group similar errors into high-level categories. As shown in the first two columns of Table 1, we classify the error patterns into three main categories: (A) Compilation Errors, (B) Execution Failures, and (C) Poor Coverage. The LLM-generated test cases exhibit various compilation errors, including import errors, namespace issues, and access errors. Most execution failures stem from incorrectly generated assertions and mocking code. Additionally, we identify poor coverage as a significant issue, primarily due to the LLMs’ inability to ensure coverage of all conditional branches and the incorrect use of assertion types.
  • Solution Guidelines: As listed in the third column of Table 1, we systematically summarize solution guidelines for the corresponding error patterns based on empirical observations of the failed test cases generated by the selected LLMs. As discussed in Section 2.3, Failed Test Case ➌ can be addressed by leveraging additional cross-file data dependencies along with Guideline (B.2). Specifically, we organize Guideline (B.2) as instructions to prompt LLMs in a knowledge-driven manner for tackling the incorrect mocking issue. In doing so, CITYWALK incorporates language-specific knowledge about mocking in C++, including best practices for creating mock objects for virtual functions. In the case of the decode method, CITYWALK provides LLMs with detailed guidance on how to correctly mock non-virtual member functions in C++. By enhancing the LLM’s understanding of these language-specific nuances, CITYWALK ensures the generated mocking code is both syntactically and semantically correct, effectively preventing execution failures caused by improper mocking of non-virtual methods. While these guidelines do not cover all error types, they are instrumental in guiding LLMs to produce high-quality test cases and mitigate common errors in C++ unit test generation.
Table 1. Language-Specific Domain Knowledge Derived from Empirical Observations
Category Error Pattern Guideline
(A) Compilation Error Import Error (A.1) Import all necessary dependencies with the correct paths.
(A.2) Use only the C++ standard libraries, imported third-party libraries, and provided methods.
(A.3) If gtest is not allowed, directly call test methods from the main function.
Namespace Error (A.4) Use the correct namespace throughout the tests.
Access Error (A.5) Properly handle static members by accessing them using the class name.
(A.6) Avoid invoking private methods or accessing private fields defined in the program.
(B) Execution Failure Incorrect Assertion (B.1) Choose appropriate assertions for the pointer data type, clearly distinguishing between
address and content comparisons.
Incorrect Mocking (B.2) For mocking (if using gmock), remember that only virtual methods can be mocked.
(C) Poor Coverage Uncovered Branch (C.1) Ensure coverage of true and false branches for each conditional predicate at least once.
Assertions with ASSERT_* (C.2) Utilize non-terminating assertions (e.g., EXPECT_*) to maximize code coverage.
2.6. Unit Test Generation
Unit test generation is a complex task that poses significant challenges when generating high-quality test cases from scratch with limited guidance. To address this, CITYWALK initially produces a test suite for the given focal method by following step-by-step instructions. Subsequently, CITYWALK corrects the failed test cases using a three-phase rule-based fixing approach. As illustrated in Algorithm 1, we provide detailed overview of the unit test generation process as follows.
Input: The given focal method: $\mathtt{m_{focal}}$; The focal contexts: $\mathtt{Context_{focal}}$; The configuration dependencies: ${\mathtt{Dep_{c}}}$; The cross-file data dependencies: $\mathtt{Dep_{d}}$; The intention contexts: $\mathtt{Context_{intent}}$; The guidelines of language-specific domain knowledge: $\mathtt{Guideline_{DK}}$; The step-by-step instruction prompt: $\mathtt{PROMPT_{step}}$; The syntactic error fixing rules: $\mathtt{Rule_{F_{s}}}$; The compilation error fixing rules: $\mathtt{Rule_{F_{c}}}$; The error fixing prompt: $\mathtt{PROMPT_{F}}$
Output: The generated unit test case(s): $\mathtt{TC}$
1 /* Initial Unit Test Case Generation via Step-by-Step Instructions */ /* Step 1: Program Understanding */
[…]
6if $\mathbf{Compiler}(\mathtt{TC})$ is not $\mathbf{PASS}$ then
7      $\mathtt{TC}\leftarrow\mathtt{Rule_{F_{c}}(TC)}$
8       if $\mathbf{Compiler}(\mathtt{TC})$ is not $\mathbf{PASS}$ then
9            $\mathtt{Context_{error}\leftarrow\mathbf{Compiler}(\mathtt{TC})}$
10            $\mathtt{TC}\leftarrow\mathbf{LLM}$($\mathtt{PROMPT_{F}(TC,Context_{error})}$
return $\mathtt{TC}$
Algorithm 1 Unit Test Case Generation for a Focal Method.
Figure 5. The Detailed Prompting Content for Generating Initial Test Cases for a Given Focal Method.
[…]
2.6.1. Initial Unit Test Case Generation via Step-by-Step Instructions
As shown in Figure 5, the prompting content for querying the LLM comprises four components: the task definition, the step-by-step instructions, the contextual information, and the output format. Within the context of CITYWALK, the LLM is employed to generate initial test cases by following steps:
  1. (1)
    Program Understanding (Line 3): First, the LLM analyzes the provided source code of $\mathtt{m_{focal}}$ to grasp its intended functionality, denoted as $\mathtt{Intent}$. Next, the LLM extracts the key elements necessary for generating effective test cases, creating a set of candidate keywords (i.e., $\mathtt{Dep_{ingredient}}$) that represent the core dependent ingredients of $\mathtt{m_{focal}}$.
  2. (2)
    Unit Test Generation (Line 5): Based on the $\mathtt{Intent}$ of $\mathtt{m_{focal}}$, the LLM generates an initial test suite for $\mathtt{m_{focal}}$ by utilizing the provided contextual information, which includes the focal contexts $\mathtt{Context_{focal}}$, project dependencies $\mathtt{Dep_{c}}$ and $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$. Additionally, the LLM matches the contextual information with the extracted $\mathtt{Dep_{ingredient}}$ to ensure an precise understanding of the purpose of $\mathtt{m_{focal}}$.
  3. (3)
    Test Case Refinement (Line 7): To enhance the effectiveness of the LLM-generated test cases $\mathtt{TC}$, the LLM refines the initial test suite using the language-specific domain knowledge as guidelines $\mathtt{Guideline_{DK}}$, ensuring the generation of high coverage and executable $\mathtt{TC}$.
Figure 6. The Three-Phase Fixing Process for the LLM-Generated Unit Test Cases.
[…]
2.6.2. Post-Processing of the LLM-Generated Unit Test Cases
Inspired by ChatUniTest (Chen et al., 2024a), CITYWALK applies post-processing techniques to fix $\mathtt{TC}$ that contain syntactic and compilation errors. As shown in Figure 6, the fixing process consists of the following three phases:
  1. Rule-based Syntactic Error Fixing: CITYWALK first utilizes $\mathtt{Rule_{F_{s}}}$ to address the syntactic errors in $\mathtt{TC}$ (Line 9). The syntactic error fixing rules include:
    • Ensuring all brackets are closed through pattern matching.
    • Removing all incorrect import statements by consolidating imports within $\mathtt{Context_{focal}}$.
    • If the project uses gtest, retaining only one main function in the generated test suite; otherwise, ensuring each test method is called by the main function.
  2. Rule-based Compilation Error Fixing: Following Phase ➊, CITYWALK compiles the generated $\mathtt{TC}$ and applies $\mathtt{Rule_{F_{c}}}$ to the failed cases (Lines 10-11). The fixing rules include:
    • Fixing incorrect usage of namespaces.
    • Deleting non-existent import paths.
  3. LLM-based Compilation Error Fixing: If compilation errors persist after the previous phase, CITYWALK transitions to a one-round LLM-based fixing phase (Lines 12-14). During this phase, CITYWALK collects details about the failed test cases along with their associated compilation error messages to construct the fixing prompt. The LLM is then prompted to analyze the root cause of the errors and make corrections to the failed test cases. If any test case continues to fail compilation after this phase, it will be removed from $\mathtt{TC}$.
3. Experimental Setup
3.1. Research Questions
To assess the effectiveness of CITYWALK, we raise the following three research questions (RQs):
  • RQ1: How does CITYWALK perform in C++ unit test generation when compared to state-of-the-art baselines? This RQ aims to evaluate the superior effectiveness of CITYWALK in comparison to open-source code LLMs, closed-source general-purpose LLMs, and LLM-based unit test generation approaches within the context of C++ unit test generation. To achieve this, we conduct a comprehensive evaluation of CITYWALK against six baselines using a diverse collection of C++ projects.
  • RQ2: How does each component impact the performance of CITYWALK? Since CITYWALK introduces a step-by-step prompting strategy and leverages various contextual information and post-processing techniques as guidance, this RQ seeks to analyze the contributions of each component through an ablation study.
  • RQ3: What is the generalizability of CITYWALK to additional LLMs? This RQ assesses the generalization capabilities of CITYWALK by extending it to two open-source LLMs tailored to the code generation task, thereby enhancing evaluation diversity.
Table 2. Statistics of the Collected Open-Source C++ Projects
Project Application Type GitHub Stars Size (LoC) # of Files # of Focal Methods # of Complex Methods Trained?
hjson-cpp  [^4] https://github.com/hjson/hjson-cpp User Interface for JSON 70 2911 4 25 9 (36.0%) Y
tinyxml2  [^5] https://github.com/leethomason/tinyxml2 XML Parser 5.2K 3606 1 158 5 (3.2%) Y
yaml-cpp  [^6] https://github.com/jbeder/yaml-cpp YAML Parser and Emitter 5.3K 8800 28 204 31 (15.2%) Y
re2  [^7] https://github.com/google/re2 Regular Expression Engine 9.1K 20373 10 146 41 (28.1%) Y
ninja  [^8] https://github.com/ninja-build/ninja Build System 11.5K 37512 19 238 38 (16.0%) Y
leveldb  [^9] https://github.com/google/leveldb Key-Value Storage Library 37K 149371 17 220 6 (2.7%) Y
json.cpp  [^10] https://github.com/jart/json.cpp JSON Parsing Library 689 62677 1 34 4 (11.8%) N
glomap  [^11] https://github.com/colmap/glomap Map Management Library 1.6K 8477 9 32 4 (12.5%) N
3.2. Benchmark
To comprehensively evaluate the quality of LLM-generated test cases, we construct a new benchmark consisting of eight real-world open-source C++ projects from GitHub. As shown in Table 2, the first four projects are widely utilized in recent studies on automated C++ unit testing (Herlim et al., 2022; Rho et al., 2023, 2024). Additionally, we crawl two practical projects related to basic software: ninja (build system) and leveldb (database). To address potential data leakage concerns, we include two projects—json.cpp and glomap—which were created after the knowledge cutoff date (October 2023). This ensures that existing LLMs have not been trained on these projects. The selection criteria for these projects are as follows. First, each project has received more than 50 stars on GitHub and is actively maintained, ensuring ongoing community interest and updates. Second, the projects span a range of application domains, including parsers (e.g., XML or YAML), basic software systems, and algorithm libraries. Third, the projects feature a number of complex methods with a cyclomatic complexity (McCabe, 1976) greater than 10, which suggests the presence of nested control flows, while also covering intricate C++ language features. The size of the selected projects varies from 2.9K lines of code (LoC) to 149.4K LoC. Larger projects are excluded to effectively manage token costs within our limited budget. In total, CITYWALK generates C++ unit test cases for 1057 focal methods across these eight projects. These selection criteria ensure that our benchmark is both high-quality and reproducible, while maintaining a balance between diversity and resource constraints.
3.3. Baselines
This paper focuses on addressing the C++ unit test generation task using LLMs. To this end, we compare CITYWALK against six state-of-the-art baselines. First, we select two representative open-source code LLMs: CodeGeeX4 (Zheng et al., 2023) and DeepSeek-V2.5 (DeepSeek-AI, 2024b, a), both of which have demonstrated competitive performance on recent public benchmarks related to coding tasks, such as BigCodeBench (Zhuo et al., 2024) and NaturalCodeBench (Zhang et al., 2024b). The selection of these two models is motivated by their distinct characteristics in terms of parameter size (CodeGeeX4 has 9B parameters, while DeepSeek-V2.5 comprises 236B parameters) and architecture (CodeGeeX4 utilizes a Transformer-based architecture, whereas DeepSeek-V2.5 employs a Mixture of Experts-based architecture). We exclude CodeLLaMA (Rozière et al., 2023) from our evaluation due to its limited context length. Additionally, we include the closed-source general-purpose LLMs (i.e., ChatGPT (OpenAI, 2022) and GPT-4o (OpenAI, 2023)) because of their established effectiveness across a broad spectrum of tasks. Finally, we compare CITYWALK with two LLM-based unit test generation approaches for Java: ChatTester (Yuan et al., 2024) and HITS (Wang et al., 2024b). Since existing automated C++ unit test generation tools, such as Coyote (Rho et al., 2023) and CITRUS (Herlim et al., 2022), are currently unavailable, we are unable to reproduce the results reported in their respective papers. Therefore, we do not include them as baselines in this paper.
3.4. Metrics
As illustrated in Algorithm 1, CITYWALK generates a test file for one focal method at a time. It executes the generated files individually and utilizes llvm-cov to compute coverage for each focal method. Specifically, this paper employs four evaluation metrics commonly used in existing studies (Yuan et al., 2024; Chen et al., 2024a; Yang et al., 2024) to compare the performance of CITYWALK with the baselines. These metrics collectively provide a comprehensive perspective on the quality, completeness, and effectiveness of the LLM-generated test cases.
  • Compilation Success Rate (CSR): This metric represents the percentage of LLM-generated test cases that compile successfully relative to the total test case number.
  • Execution Pass Rate (EPR): This metric quantifies the percentage of LLM-generated test cases that pass during execution, reflecting the proportion of tests that yield expected outcomes.
  • Line Coverage ($\mathbf{Cov_{L}}$): This metric assesses the percentage of source code lines within the focal methods that are executed by LLM-generated test cases.
  • Branch Coverage ($\mathbf{Cov_{B}}$): This metric evaluates the percentage of logical conditions in the source code that are explored by LLM-generated test cases.
3.5. Implementation
We implement the core logic of CITYWALK in Python, invoking GPT-4o via its API, specifically using the gpt-4o version from the GPT family of models, which is recognized as the most advanced model currently available. We do not truncate input prompts, as the selected LLMs can effectively handle lengthy inputs. We set a maximum output limit of 4096 tokens. In all experiments, we utilize greedy decoding to generate responses, with CITYWALK producing the top-1 chat completion choice for each input prompt. To enhance response stability, we set the sampling temperature to 0. Additionally, we conduct experiments in a zero-shot setting, where no task examples are provided, thereby demonstrating the superiority of CITYWALK.
4. Results and Analysis
4.1. Answering RQ1
To answer this question, we conduct a comprehensive comparison of CITYWALK against six state-of-the-art baselines using the collected benchmark. For each LLM baseline, we utilize its inference API for implementation, employing the same basic prompt described in Section 1. For ChatTester (Yuan et al., 2024) and HITS (Wang et al., 2024b), we use their open-source reproduction artifacts for implementation. Furthermore, we apply the same configuration settings as those used in CITYWALK for fair comparison.
Table 3. Comparison of CITYWALK against the Baselines in Terms of Compilation Success Rate (CSR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 13.56%
DeepSeek-V2.5 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 14.47%
ChatGPT 0.00% 0.00% 19.34% 5.45% 5.00% 9.64% 31.16% 28.57% 12.40%
GPT-4o 16.17% 41.26% 22.01% 0.00% 6.90% 9.74% 51.62% 49.70% 24.68%
ChatTester 20.00% 18.02% 6.42% 4.68% 11.36% 0.00% 18.75% 8.82% 11.01%
HITS 1.10% 5.94% 5.65% 0.00% 20.37% 2.66% 0.00% 1.92% 4.71%
CITYWALK 91.84% 97.25% 65.01% 64.85% 47.34% 47.49% 100.00% 47.65% 70.18%
4.1.1. Experimental Metric Evaluation
Table 3, Table 4, Table 5, and Table 6 present the performance of CITYWALK and selected baselines in C++ unit test generation, evaluated across both correctness metrics (i.e., CSR and EPR) and coverage metrics (i.e., $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$). The best result for each metric is highlighted in bold, while the second-best result is underlined. Our experiments yield the following key findings:
  1. (1)
    CITYWALK demonstrates superior performance compared to state-of-the-art baselines on the collected benchmark. When compared to the six LLM-based baselines, CITYWALK achieves the best results across all evaluation metrics (except CSR and EPR in the glomap project). A closer examination of the results reveals that the second-best outcomes for each metric vary depending on the project under test. Nevertheless, GPT-4o consistently outperforms other baselines regarding average scores of all the four evaluation metrics (listed in the Avg. rows). Specifically, CITYWALK surpasses the best baseline GPT-4o by 42.05%, 32.86%, 26.41%, and 23.09% in terms of CSR, ERP, $\mathbf{Cov_{L}}$, and $\mathbf{Cov_{B}}$, respectively. These improvements underscore the effectiveness of CITYWALK in the C++ unit test generation task.
    Table 4. Comparison of CITYWALK against the Baselines in Terms of Execution Pass Rate (EPR)
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 4.69%
    DeepSeek-V2.5 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 12.05%
    ChatGPT 0.00% 0.00% 18.37% 3.41% 4.04% 8.66% 30.82% 10.48% 9.47%
    GPT-4o 14.37% 38.46% 15.75% 0.00% 6.57% 7.61% 49.10% 47.88% 22.47%
    ChatTester 12.00% 13.95% 4.91% 1.75% 8.77% 0.00% 9.38% 8.82% 7.45%
    HITS 1.10% 5.45% 4.52% 0.00% 14.90% 1.73% 0.00% 1.92% 3.70%
    CITYWALK 57.14% 88.24% 51.03% 49.37% 42.47% 41.04% 98.91% 38.26% 58.31%
    Table 5. Comparison of CITYWALK against the Baselines in Terms of Line Coverage ($\mathbf{Cov_{L}}$)
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 9.95%
    DeepSeek-V2.5 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 17.75%
    ChatGPT 0.00% 0.00% 25.59% 0.84% 10.68% 14.73% 12.09% 16.75% 10.09%
    GPT-4o 7.73% 43.21% 30.17% 0.00% 10.64% 15.45% 26.50% 22.31% 19.50%
    ChatTester 1.68% 45.00% 0.00% 3.05% 17.52% 0.00% 3.56% 4.47% 9.41%
    HITS 0.00% 35.64% 25.56% 0.00% 19.55% 8.53% 0.00% 4.15% 11.68%
    CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15% 45.91%
    Table 6. Comparison of CITYWALK against the Baselines in Terms of Branch Coverage ($\mathbf{Cov_{B}}$)
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 9.47%
    DeepSeek-V2.5 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 16.00%
    ChatGPT 0.00% 0.00% 21.61% 0.51% 10.22% 12.50% 13.06% 13.33% 8.90%
    GPT-4o 8.82% 37.18% 23.54% 0.00% 8.97% 14.06% 24.84% 17.78% 16.90%
    ChatTester 0.89% 38.45% 0.00% 2.24% 15.46% 0.00% 1.61% 2.48% 7.64%
    HITS 0.00% 26.74% 17.70% 0.00% 16.46% 8.72% 0.00% 3.09% 9.09%
    CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82% 39.99%
  2. (2)
    CITYWALK effectively generates fewer compilation errors than the selected baselines. As observed from Table 3, it is important to that, in many instances, the selected baselines fail to generate any compilable unit test cases for the projects under test using the basic prompt. This issue arises because projects such as hjson-cpp, tinyxml2, re2, and json.cpp do not include the gtest testing framework. Consequently, when the LLMs generate test code without accounting for these configuration dependencies, they default to using the gtest framework, resulting in compilation errors for all generated test cases. In contrast, the higher CSR achieved by CITYWALK is largely due to the inclusion of additional project dependencies. CITYWALK guides the LLMs to generate syntactically correct test code by directly incorporating relevant configuration dependency information into the prompt, thereby greatly reducing the occurrence of compilation errors.
    Table 7. Comparison of CITYWALK against the Baselines in Terms of the Number and Complexity of Generated Test Cases
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. Num Avg. Size Avg. CC
    CodeGeeX4 175 417 652 1133 1231 1115 402 193 664.8 7.4 1.1
    DeepSeek-V2.5 463 1405 1147 1466 2273 1572 533 233 1136.5 7.5 1.1
    ChatGPT 59 172 827 587 1139 1235 292 105 552.0 6.8 1.1
    GPT-4o 167 143 750 777 1218 986 277 165 560.4 8.1 1.2
    ChatTester 25 172 265 171 308 163 32 34 146.3 7.1 1.3
    HITS 181 404 1239 1190 1389 752 233 156 693.0 6.0 1.0
    CITYWALK 49 255 243 219 657 386 183 123 264.4 9.7 1.5
  3. (3)
    The project’s complexity significantly affects the ability of CITYWALK to generate correct test cases with high coverage. According to the statistical results in the seventh column of Table 2, tinyxml2 exhibits a smaller proportion of complex methods. Consequently, the test cases generated by CITYWALK for tinyxml2 achieve both high syntactical correctness, with a CSR exceeding 95%, and successful execution, with an EPR above 85%. This leads to strong coverage metrics, with both $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$ exceeding 65%. Conversely, for the more complex project re2, which contains the most complex methods, the correctness of the generated test code is comparatively low, and the coverage metrics are also less satisfactory. Additionally, although leveldb has relatively few complex methods, CITYWALK performs poorly in generating test cases for this project. This underperformance can be attributed to several factors. We will discuss the intricate cases from leveldb in Section 4.1.2.
  4. (4)
    CITYWALK achieves high code coverage with fewer test cases than the baselines. Table 7 reports the number of test cases generated by CITYWALK and the baselines for each project. The Avg. Num represents the average number of generated test cases across all evaluated projects, while the Avg. Size and Avg. CC represent the average source code lines and cyclomatic complexity of all generated test cases, respectively. The statistical results indicate that CITYWALK generates significantly fewer test cases compared to the selected baselines (except ChatTester), proving that CITYWALK does not rely on sampling a large number of test cases to achieve advantages. The reason ChatTester generates the fewest test cases lies in its prompt design, which explicitly instructs the LLM to “write one test case for each focal method”. This constraint significantly limits the number of test cases generated by ChatTester, resulting in suboptimal performance of ChatTester across many projects. Meanwhile, the scale and complexity of the test cases generated by CITYWALK are slightly higher than those of the LLMs. This outcome is reasonable, as generating correct test assertions necessitates a certain level of complex code logic.
  5. (5)
    The correctness and coverage metrics do not necessarily indicate a positive correlation. As illustrated in Tables 3, Tables 4, Tables 5 and Tables 6, high correctness does not always guarantee high code coverage, and low correctness can sometimes lead to higher coverage. For instance, DeepSeek-V2.5 on yaml-cpp demonstrates this phenomenon. Despite achieving lower CSR and EPR compared to other LLM baselines, DeepSeek-V2.5 attains the second-highest code coverage scores. According to the fourth column in Table 7, it is evident that DeepSeek-V2.5 generates more test cases for yaml-cpp than other LLM baselines. Consequently, the proportion of incorrect test cases is also relatively high, leading to a scenario characterized by low correctness but high coverage.
  6. (6)
    The two LLM-based unit test generation approaches for Java (i.e., ChatTester and HITS) achieve low performance on the collected C++ projects. First, the output of LLMs can vary significantly depending on the prompt design. If the prompt does not account for the specific characteristics of different programming languages, the performance of existing Java-specific baselines on C++ unit test generation may not generalize effectively. Furthermore, to ensure fairness in our experiments, all baselines that use LLMs for post-processing perform only a single round of iterative fixes, similar to CITYWALK, which is much fewer than the number of iterations used in their original papers. As a result, the reproduction performance may be lower than the corresponding results reported. However, this also suggests that CITYWALK is less dependent on iterative fixes to achieve its performance.
Figure 7. Error Category and Frequency.
[…]
4.1.2. Bad Case Breakdown
We further conduct an in-depth investigation into the common error categories of the failed test cases generated by CITYWALK. Specifically, we manually analyze the output error messages associated with these failed test cases and summarize the corresponding error categories. In total, CITYWALK produces 1329 failed test cases across eight projects. Our analysis reveals the following findings based on the statistical results:
  1. (1)
    The two most prevalent categories of errors in the failed test cases are compilation errors and execution errors. As illustrated in Figure 7, we subdivide the two high-level categories into distinct sub-categories. Due to space constraints, we only present the errors that are observed more than ten times. The most common sub-category of compilation errors is Access Error, which occurs when LLMs frequently generate test code that attempts to invalidly access private variables or methods. The most common sub-category of execution errors is Test Setup Error, primarily due to LLMs facing challenges in accurately inferring the mock object configurations, as well as generating valid test data inputs.
  2. (2)
    The complexity of C++ language features poses significant challenges for CITYWALK in generating accurate test cases. In the design of CITYWALK, we integrate language-specific domain knowledge derived from empirical observations to address common errors, thereby enhancing the performance of the LLM to some extent. However, since the benchmark is collected from a diverse range of real-world projects, project-specific issues may still impede the generation of correct and high-coverage test code. For instance, many functionalities in the large basic software system leveldb rely heavily on external interactions, such as disk I/O operations and file content checks. This dependence often compels LLMs to redefine existing methods in dependent classes or to invoke undefined methods when constructing I/O streams, leading to errors such as Multiple Definition Error and Undefined Symbol Error, which result in low evaluation metrics. Additionally, leveldb involves advanced features like synchronization locks, which significantly increase the complexity and challenge the test generation process. Moreover, the absence of an automated and strict exception handling mechanism, akin to that in Java, complicates the LLM’s ability to generate correct test code with exception handling logic. The most frequently occurring assertion errors can be traced back to the lack of supplementary post-processing techniques, which could help prompt LLMs to rectify incorrect assertions in the generated test cases. We are concerned that providing LLMs with error messages that include expected outputs may encourage them to replicate those outputs directly, thus manipulating the compiler into successful execution of the generated test cases. We will discuss these false positive cases in detail in Section 5.2.
Figure 8. The Ablation Study for CITYWALK in Terms of the Coverage Metrics.
[…]
Answer to RQ1: In conclusion, CITYWALK exhibits a marked superiority over the LLM-based baselines across all evaluation metrics, highlighting its effectiveness in C++ unit test generation.
4.2. Answering RQ2
To answer this question, we conduct a series of ablation experiments to assess the impact of different designed components within CITYWALK, including configuration dependencies ${\mathtt{Dep_{c}}}$, cross-file data dependencies $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$, guidelines of language-specific domain knowledge $\mathtt{Guideline_{DK}}$, step-by-step instructions $\mathtt{PROMPT_{step}}$, and post-processing techniques $\mathtt{Fix_{rule+prompt}}$. To ensure fairness in comparisons, the parameter configurations align with those described in Section 3.5. Specifically, we perform ablation experiments by removing one component at a time and analyze the performance contribution of each component regarding the coverage metrics $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$. Figure 8 illustrates the comparison results of CITYWALK against the ablated models across all evaluated projects. When compared to CITYWALK, each ablation model exhibits varying degrees of decline in terms of $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$, indicating that each designed component contributes to the overall enhancement in generating high-quality unit test cases. Table 8 and Table 9 show the performance contribution results (denoted as the performance degradation values). The greater the coverage scores decline, the larger the contribution of that component. The Avg. $\downarrow$ columns present the average results of all projects. According to the statistical results presented in Table 8 and Table 9, the top three components that contribute the most to CITYWALK are ${\mathtt{Dep_{c}}}$, ${\mathtt{Dep_{d}}}$, and $\mathtt{Context_{intent}}$. As discussed in Section 4.1.1, ${\mathtt{Dep_{c}}}$ plays a crucial role in guiding LLMs to generate syntactically correct test code by directly incorporating relevant configuration dependencies. Therefore, projects such as hjson-cpp, tinyxml2, re2, and json.cpp, which do not use gtest, would benefit from ${\mathtt{Dep_{c}}}$ by helping the LLMs understand the absence of such a testing framework, thereby preventing the generation of non-compilable test cases. Figure 9 illustrates how extracted cross-file data dependencies are used as guidance for LLMs to generate correct test cases. By supplementing the key $\mathtt{Dep_{d}}$ (i.e., the initialization constructor of Json) from the json.h file, LLMs can avoid errors that would otherwise arise from using non-existent functions (i.e., setNumber). Meanwhile, Figure 10 illustrates how $\mathtt{Context_{intent}}$ can guide LLMs in generating correct test cases for the Failed Test Case ➋ within Figure 1. Specifically, the retrieved ParseTag method within the singledocparser class provides an invocation example of the focal method, while the retrieved _Tag method within the emittermanip header file offers an initialization example of the focal method. These examples significantly aid LLMs in understanding the usage patterns of the focal method, contributing to the generation of functionally correct assertions and enhancing code coverage.
Table 8. Performance Contribution of Each Component in Terms of Line Coverage ($\mathbf{Cov_{L}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\downarrow$
CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15%
   w/o ${\mathtt{Dep_{c}}}$27.94% -77.31% -12.08% -42.56% -3.82% -1.36% -50.71% -1.81% -27.20%
   w/o ${\mathtt{Dep_{d}}}$0.82% -32.42% -14.50% -7.07% -15.26% -8.69% -19.04% -13.03% -13.85%
   w/o ${\mathtt{Context_{intent}}}$12.36% -29.84% -8.72% -8.21% -11.21% -5.11% -6.59% -9.70% -11.47%
   w/o ${\mathtt{Guideline_{DK}}}$16.74% -7.40% -11.48% -7.87% -7.04% -2.69% -3.59% -5.59% -7.80%
   w/o ${\mathtt{PROMPT_{step}}}$5.46% -12.97% -12.33% -9.86% -8.42% -8.91% -12.48% -7.59% -9.75%
   w/o ${\mathtt{Fix_{rule+prompt}}}$5.24% -12.51% -14.95% -8.48% -10.76% -1.51% -1.27% -4.26% -7.37%
Table 9. Performance Contribution of Each Component in Terms of Branch Coverage ($\mathbf{Cov_{B}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\downarrow$
CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82%
   w/o ${\mathtt{Dep_{c}}}$24.29% -65.66% -10.04% -35.64% -3.38% -1.50% -46.61% -1.70% -23.60%
   w/o ${\mathtt{Dep_{d}}}$1.99% -30.69% -11.48% -7.94% -14.60% -8.75% -19.33% -13.93% -13.59%
   w/o ${\mathtt{Context_{intent}}}$14.56% -28.48% -7.00% -9.87% -9.60% -5.17% -6.83% -9.59% -11.39%
   w/o ${\mathtt{Guideline_{DK}}}$15.51% -8.07% -8.70% -9.02% -5.83% -2.75% -3.72% -5.70% -7.41%
   w/o ${\mathtt{PROMPT_{step}}}$5.97% -13.60% -10.19% -9.90% -6.16% -8.97% -12.49% -7.59% -9.36%
   w/o ${\mathtt{Fix_{rule+prompt}}}$6.12% -9.68% -8.52% -9.60% -$3$11% -0.64% -3.26% -8.91%
Figure 9. Illustration Example of How Cross-File Data Dependencies Guide LLMs Generate Correct Test Cases.
[…]
Figure 10. Illustration Example of How Intention Contexts Guide LLMs Generate Correct Test Cases.
[…]
Answer to RQ2: To sum up, all components of CITYWALK significantly improve the performance of C++ unit test generation in terms of the coverage metrics.
4.3. Answering RQ3
To answer this question, We further employ CITYWALK with two open-source code LLMs, namely CodeGeeX4-9B and DeepSeek-V2.5-236B. Specifically, we conduct ablation experiments on each LLM individually to investigate the impact of querying the corresponding LLM using the prompting strategy and additional contextual guidance designed by CITYWALK. Table 10, Table 11, Table 12 and Table 13 present the comparison results using the correctness and coverage metrics, respectively. Each LLM’s results are displayed in two lines: the first line shows the results when the LLM directly utilizes the basic prompt to generate unit test cases for the given focal methods, and the second line presents the results when integrated with CITYWALK. We derive two key insights from the statistical results within the tables: Integrating CITYWALK as a complementary plug-in enhances the performance of C++ unit test generation. As observed, both code LLMs demonstrate consistent improvements in all four evaluation metrics across the eight projects. The performance gains exhibit a marked increase with the number of parameters. The average performance gain of each metric (listed in the Avg. $\uparrow$ column) with DeepSeek-V2.5-236B is substantially greater than that of CodeGeeX4-9B.
Table 10. Generalization Results of CITYWALK on Different Code LLMs in Terms of Compilation Success Rate (CSR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 21.96%
CodeGeeX4-9B w/ CITYWALK 43.59% 42.11% 54.64% 4.30% 31.45% 29.12% 36.45% 42.44%
DeepSeek-V2.5 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 40.39%
DeepSeek-V2.5-236B w/ CITYWALK 74.62% 72.67% 43.45% 47.83% 41.56% 42.53% 53.89% 62.34%
Table 11. Generalization Results of CITYWALK on Different Code LLMs in Terms of Execution Pass Rate (EPR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 22.84%
CodeGeeX4-9B w/ CITYWALK 30.77% 37.92% 42.92% 2.15% 26.78% 25.56% 21.78% 32.34%
DeepSeek-V2.5 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 35.46%
DeepSeek-V2.5-236B w/ CITYWALK 67.69% 64.92% 31.82% 42.61% 36.78% 37.63% 47.12% 51.56%
Table 12. Generalization Results of CITYWALK on Different Code LLMs in Terms of Line Coverage ($\mathbf{Cov_{L}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 17.75%
CodeGeeX4-9B w/ CITYWALK 15.78% 46.15% 35.28% 28.04% 19.12% 12.45% 26.89% 37.89%
DeepSeek-V2.5 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 26.47%
DeepSeek-V2.5-236B w/ CITYWALK 22.44% 78.17% 42.80% 39.71% 52.23% 26.63% 48.35% 43.45%
Table 13. Generalization Results of CITYWALK on Different Code LLMs in Terms of Branch Coverage ($\mathbf{Cov_{B}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 12.68%
CodeGeeX4-9B w/ CITYWALK 5.18% 33.07% 31.23% 21.09% 21.45% 10.72% 21.53% 32.98%
DeepSeek-V2.5 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 23.65%
DeepSeek-V2.5-236B w/ CITYWALK 23.65% 65.98% 35.06% 31.94% 51.89% 25.15% 45.41% 38.12%
Answer to RQ3: CITYWALK effectively harnesses the latent intelligence of the LLMs, demonstrating the potential for seamless integration with additional LLMs in a plug-and-play manner.
Figure 11. Illustration of Two Real-World Issues from tinyxml2 Found by CITYWALK.
[…]
5. Discussion
5.1. Usefulness of CITYWALK in Finding Real-World Bugs
Finding real software bugs is one of the most important criteria for evaluating an automated unit test generation approach. Therefore, we conduct a case-by-case analysis to assess the ability of CITYWALK to trigger exceptions and discover bugs. First, we select two bugs from tinyxml2 that were identified by CITYWALK-generated test cases. These are real-world issues that have been reported by human developers. As illustrated in Figure 11(a), the CITYWALK-generated test case first creates an XMLDocument object, doc, which then invokes the Value function. This function, in turn, calls XMLNode::Value(). Since the variable _value within XMLNode::Value() may not have been initialized, it triggers an assertion error in the GetStr function. This bug, identified by CITYWALK, corresponds to Issue #323  [^12] https://github.com/leethomason/tinyxml2/issues/323 in tinyxml2, where a human developer also commits an assertion failure when calling Value. Figure 11(b) shows another bug triggered by CITYWALK, where invalid hexadecimal format string arguments are passed to ToInt64. This bug is associated with Issue #825  [^13] https://github.com/leethomason/tinyxml2/issues/825 in tinyxml2, where a human developer also identifies a similar issue when calling the static member function ToInt64. In summary, the test cases generated by CITYWALK are effective in identifying real bugs within open-source projects.
5.2. False-Positive Executable Test Cases
Existing LLM-based unit test generation approaches (Yuan et al., 2024; Wang et al., 2024b) primarily rely on compiler error messages to guide LLMs in an iterative fixing process after generated test cases fail to execute. CITYWALK excludes such a fixing strategy because the error messages inherently contain the actual outcomes of the focal methods. This can lead LLMs to adjust expected values based on actual results or manipulate outputs to match expected outcomes, thereby passing tests under misleading circumstances. As illustrated in Figure 12, when provided with the error message, GPT-4o directly uses the expected value "str" as the input for assertions, resulting in a passing test case that is, in reality, a false positive. This behavior is undesirable, as it undermines the reliability of unit testing and fails to verify the intended functionality of the focal method. Thus, effectively addressing the false positive issue remains as our future work.
Figure 12. False Positive Example for the Failed Test Case ➋ within Figure 1 using Iterative LLM-Based Fixing.
[…]
5.3. Threats to Validity
The primary threat to the validity of CITYWALK lies in the diversity of the projects used for evaluation and its generalization to other LLMs. In this paper, we collect 1057 focal methods across eight real-world open-source C++ projects crawled from GitHub, ensuring a degree of diversity and quality in our evaluation. For comparison, we select two open-source code LLMs (CodeGeeX4 and DeepSeek-V2.5) and two closed-source commercial LLMs (ChatGPT and GPT-4o), considering their varied model sizes, architectures, and effectiveness on coding tasks. Future work will involve expanding the benchmark and integrating additional LLMs to better assess the generalizability of CITYWALK.
6. Related Work
To mitigate the manual effort associated with writing unit tests for developers in practice, researchers have proposed various automation techniques aimed at enhancing testing efficiency. Existing approaches can be broadly categorized into the following three technical avenues.
6.1. Program Analysis-Based Automated Unit Testing Tools
EvoSuite (Fraser and Arcuri, 2011) is an automated test case generation tool tailored for Java. It utilizes mutation testing and constraint-solving techniques to generate appropriate assertions, which effectively summarize the behavior of the program while maximizing the number of killed mutants. In contrast, Randoop (Pacheco et al., 2007; Pacheco and Ernst, 2007) is an automated tool that adopts feedback-driven random testing techniques to generate assertions. Randoop leverages the outcomes of test executions to generate assertions that accurately capture the program’s behavior. Pynguin (Lukasczyk and Fraser, 2022), an extendable test generation framework for Python to produce regression tests via search-based techniques. Coyote C++ (Rho et al., 2023) employs a sophisticated concolic execution-based method to facilitate fully automated unit testing for C and C++. Despite demonstrating commendable performance in achieving high code coverage, the aforementioned tools exhibit certain limitations: (1) The test code generated by existing tools often suffers from poor readability (Daka et al., 2015); (2) The assertions produced by these tools are frequently insufficient in effectively detecting real-world faults (Shamshiri et al., 2015; Shamshiri, 2015); (3) Search-based techniques may encounter path explosion issues due to excessively large search spaces (Tang et al., 2024).
6.2. Pre-Trained Language Model-Based Automated Unit Test Generation
Tufano et al. (Tufano et al., 2021) pre-trained a language model on large-scale unsupervised Java corpora, subsequently fine-tuning the model for unit test generation, thereby enabling the efficient generation of test cases. Zhang et al. (Zhang et al., 2023) adopted the summarization of focal methods as complementary information to capture the developers’ intent, which aids in generating meaningful test assertions for helping developers in writing accurate unit test cases. Similarly, Alagarsamy et al. (Alagarsamy et al., 2024) utilized both focal methods and assertion declarations during model pre-training, aiming to establish connections between focal methods and corresponding test cases. Shin et al. (Shin et al., 2024) developed project-specific datasets for domain adaptation by leveraging existing developer-written test cases within each project, promoting the generation of more human-like unit tests. Steenhoek et al. (Steenhoek et al., 2023) employed reinforcement learning for model optimization, designing reward functions based on the static quality metrics of the generated unit test cases. Nonetheless, unit test cases generated through the paradigm of pre-training and fine-tuning frequently encounter compilation or execution failures (Yang et al., 2022).
6.3. LLM-based Automated Unit Test Generation
With the emergence of generative artificial intelligence, researchers have increasingly investigated LLM-driven approaches for automatically generating unit test cases (Kang et al., 2023; Wang et al., 2024a). ChatUniTest (Chen et al., 2024a) adaptively constructs dependency contexts for the focal method, and employs a generate-validate-fix mechanism to address errors in the generated test cases. Similarly, ChatTester (Yuan et al., 2024) enhances the quality of generated test cases via intent comprehension and iterative correction. While both approaches mitigate the absence of certain dependency information, they fail to account for the influence of project dependencies, such as environment requirements, on the process of test case generation, leaving opportunities for performance optimization. SymPrompt (Ryan et al., 2024), on the other hand, guides LLMs to generate high-coverage test suite code by incorporating symbolic execution-based path information into the prompts. However, SymPrompt does not address the issue of path reachability, which can result in the inclusion of unreachable path information in the prompts, ultimately leading to inaccurate model reasoning. HITS (Wang et al., 2024b) simplifies the analysis of the focal method through program decomposition, achieving high coverage scores by prompting LLMs to generate test cases for sliced code segments. In this work, our proposed framework not only performs an in-depth analysis of the project under test but also utilizes language-specific knowledge to ensure the accuracy and reliability of the unit test cases generated by LLMs.
7. Conclusion
This paper presents a novel framework CITYWALK designed to enhance the capabilities of LLMs in generating high-quality C++ unit test cases. We explore the potential of GPT-4o by integrating program analysis techniques with retrieval-augmented strategies, providing guidance through three key aspects: project dependencies, intention contexts, and language-specific knowledge. Additionally, we decompose the unit test generation task into three distinct stages, utilizing step-by-step instructions to streamline the generation of C++ test cases, complemented by effective post-processing techniques. Extensive experiments demonstrate the superiority of CITYWALK, and further ablation studies validate the contributions of each designed component within CITYWALK.